Please help,
Given the following system of equations
x + y = 4k
2x = k
where k is a constant Real Number, solve the system for (x; y) in terms of k. Do not replace the constant k with a numerical value. Give your solution as an ordered pair.

Answer by JoelSchwartz(130)

2x=k
2x/2=k/2
x=k/2
x+y=4k
k/2+y=4k
y=4k-k/2
y=(8/2)k-k(1/2)
y=(7/2)k
(k/2; 7k/2)